# MultiNash-PF: A Particle Filtering Approach for Computing Multiple Local Generalized Nash Equilibria in Trajectory Games

### v1

@@ VI Numerical Simulations
With this setup, we employ MultiNash-PF from Algorithm 1 to efficiently capture the multi-modality in the solution space and then extract the solutions to provide a warm start trajectory to a solver like IPOPT [23] to recover different local GNEs. We use the Julia programming language and utilize the JuMP [34] library to solve the game using IPOPT solver. The results of our method are provided in Fig 1. As we can see, the presence of two agents and collision avoidance constraints induces two different possible Nash equilibria. Either both the agents will yield to their right, or they will yield to their left to avoid collision. As shown in Fig 1, our method can recover both equilibria.
@@
We consider the same unicycle dynamics as before, and the reference trajectory cost is also the same as in (16). Furthermore, the collision avoidance radius between the two agents is also the same as before. In addition to (17), the constraints for this game will include both agents avoiding the obstacle $g^{i}_{o}(x_{t},u_{t}):=\begin{bmatrix}-d(x^{i}_{t},x_{\text{obs}})+r_{\text{obs}}\end{bmatrix}$ for $i=1,2$ where $x_{\text{obs}}$ is the location of the obstacle. We also assume that we have constraints on the control bounds $g_{b}(x_{t},u_{t})=\begin{bmatrix}-v^{1}_{t},&-v^{2}_{t},&|u^{1}_{t}|-u^{1}_{b},&|u^{2}_{t}|-u^{2}_{b}\end{bmatrix}$, where $u^{1}_{b}$ and $u^{2}_{b}$ are controlled action bounds for agents 1 and 2, respectively. We write the combined constraints as $g(x_{t},u_{t}):=\begin{bmatrix}g_{c}&g^{1}_{o}&g^{2}_{o}&g_{b}\end{bmatrix}^{\top}\leq 0.$ We choose the values of control action bounds to be $u^{1}_{b}=u^{2}_{b}=[0.15,0.75]^{\top}$. This setup induces multitudes of local equilibria. Two trivial local equilibria are when the two agents pass each other through opposite sides of the obstacle. However, there exist four other nontrivial equilibria when both agents can go on the same side of the obstacle, and one of the agents can yield to the other one. It is challenging to recover all six equilibria without utilizing the capabilities of the implicit particle filter. Upon performing simulation experiments with our method, we observe that MultiNash-PF is able to recover all six Nash equilibria as shown in Fig. 2.
@@
Figure 3: Histograms of the comparison of MultiNash-PF with the baseline for the experiment shown in Fig. 2.
@@
VI-1 Comparison with Baseline
In order to compare the effectiveness of our method, we consider a baseline method. In the baseline method, we directly initialize the IPOPT solver with random perturbations of reference trajectories and compute the resulting optimal solutions. For our method, we fixed $J=50$ particles and obtained the solution trajectories using implicit particle filtering from Algorithm 1 described in Section V. After that, we extract all the different equilibria and provide them as a warm starting point for IPOPT to obtain the final solutions. The compute time of our method is a combination of time used in implicit particle filtering and time used by IPOPT to obtain all the different equilibria given warm starting from filtering.
@@ VII Conclusion
In this work, we introduced MultiNash-PF, a novel algorithm designed to efficiently compute multiple local generalized Nash equilibria (GNEs) in constrained trajectory games. By integrating potential game theory with implicit particle filtering, MultiNash-PF provides coarse estimates of local minimizers, which are then refined using off-the-shelf optimization solvers. Our experimental results in two-player trajectory games involving nonlinear dynamics and collision avoidance constraints demonstrate that MultiNash-PF not only identifies all local Nash equilibria but also significantly reduces computation time compared to traditional random initialization methods. This enhanced efficiency and accuracy underline the potential of MultiNash-PF to improve identifying various interaction modes in complex, real-world environments, paving the way for more robust and socially acceptable autonomous systems.

### v2

@@ V Numerical Simulations
With this setup, we employ MultiNash-PF from Algorithm 1 to efficiently capture the multi-modality in the solution space and then extract the solutions using hierarchical clustering [35] to provide a warm start trajectory to IPOPT [24] solver to recover different local GNEs. We use the Julia programming language and utilize the JuMP [38] library to solve the game using the IPOPT solver. The results of our method are provided in Fig 2. As we can see, the presence of two agents and collision avoidance constraints induces two different possible GNEs. Either both the agents will yield to their right, or they will yield to their left to avoid collision. As shown in Fig 2, our method can recover both equilibria.
@@
We consider the same unicycle dynamics as before, and the reference trajectory cost is also the same as in (13). Furthermore, the collision avoidance radius between the two agents is also the same as before. In addition to (14), the constraints for this game will include both agents avoiding the obstacle $g^{i}_{o}(x_{t},u_{t}):=\begin{bmatrix}-d(x^{i}_{t},x_{\text{obs}})+r_{\text{obs}}\end{bmatrix}$ for $i=1,2$ where $x_{\text{obs}}$ is the location of the obstacle. We also assume that we have constraints on the control bounds $g_{b}(x_{t},u_{t})=\begin{bmatrix}-v^{1}_{t},&-v^{2}_{t},&|u^{1}_{t}|-u^{1}_{b},&|u^{2}_{t}|-u^{2}_{b}\end{bmatrix}$, where $u^{1}_{b}$ and $u^{2}_{b}$ are controlled action bounds for agents 1 and 2, respectively. We write the combined constraints as $g(x_{t},u_{t}):=\begin{bmatrix}g_{c}&g^{1}_{o}&g^{2}_{o}&g_{b}\end{bmatrix}^{\top}\leq 0.$ We choose the values of control action bounds to be $u^{1}_{b}=u^{2}_{b}=[0.15,0.75]^{\top}$. This setup induces multitudes of local equilibria. Two trivial local equilibria are when the two agents pass each other through opposite sides of the obstacle. However, there exist four other nontrivial equilibria when both agents can go on the same side of the obstacle, and one of the agents can yield to the other one. It is challenging to recover all six equilibria without utilizing the capabilities of the implicit particle filter. Upon performing simulation experiments with our method, we observe that MultiNash-PF can recover all six Nash equilibria as shown in Fig. 4 and Fig. 3.
@@
Figure 5: Comparison of MultiNash-PF with the baseline for the experiment shown in Fig. 3.
@@
V-1 Comparison with Baseline
In order to compare the effectiveness of our method, we consider a baseline method. In the baseline method, we directly initialize the IPOPT solver with random perturbations of reference trajectories and compute the resulting optimal solutions. For our method, we fixed $J=50$ particles and obtained the solution trajectories using implicit particle filtering from Algorithm 1 described in Section IV. After that, we extract all the different equilibria and provide them as a warm starting point for IPOPT to obtain the final solutions. The compute time of our method is a combination of time used in implicit particle filtering and time used by IPOPT to obtain all the different equilibria given warm starting from filtering.
@@ V-1 Comparison with Baseline
VI Hardware Experiment
Finally, we conduct a hardware experiment where a human and a robot exchange their position as shown in Fig. 1. We perform these experiments using a TurtleBot 4 which is a ground robot with unicycle dynamics. We assume that the human also follows unicycle dynamics. The collision avoidance constraint between the TurtleBot and human is assumed to be the same as of (14). We use the Vicon motion capture system to measure the current position of both the robot and the human.
Similar as shown in Fig. 2, two GNEs may exist which will correspond to two different modes of interactions. Namely, either the human and robot yield to their right or they yield to their left. For the robot to safely navigate around the human without collision, the robot needs to be aware of both the GNEs and needs to be able to determine which GNE the human chooses. Using MultiNashPF, the robot first uses an implicit particle filter to obtain the coarse estimates of the two different equilibria and then utilizes them as a warm start for IPOPT to obtain two modes of interaction. We assume that the interaction happens over $T=10s$. Then similar to [11], the robot uses a potential function based game theoretic planner in a model-predictive fashion with a planning horizon of 5s and $\Delta t=0.1s$ to plan its motion.
The robot faces two choices of GNEs to follow recovered from the MultiNashPF. At any given time $t$ of planning, the robot computes the Fréchet distance [37] between the trajectory exhibited by the human so far and the human’s trajectory in both the computed GNEs. Let the distances of human trajectory from both the GNEs be denoted by $d_{1}$ and $d_{2}$, respectively. The robot keeps an uncertainty on the GNE that the human chooses until $|d_{1}-d_{2}|>d^{*}$ where $d^{*}$ is the specified threshold. For example, in Fig. 1 at $t=2$, values of $d_{1}$ and $d_{2}$ are almost identical which indicates that human has not chosen their mode of interaction. Whenever $|d_{1}-d_{2}|>d^{*}$ occurs, the robot becomes certain of the mode of interaction followed by the human. Then the robot follows GNE corresponding to $\arg\min_{i\in\{1,2\}}\{d_{i}\}$. For example, in Fig. 1, at $t=4$, the Fréchet distance of the human from its first GNE trajectory is very high while for the second GNE, the Fréchet distance is low indicating that the human has chosen the second GNE which corresponds to both the agents yielding to their right. Using our method, the robot autonomously identifies this mode, moves accordingly, and reaches its goal location while successfully avoiding collision with the human.
We ran 10 different trials and each time we asked the human to either yield to their right or left randomly. Our algorithm was able to successfully compute both modes and identify the mode that the human was following. Using our algorithm, the robot was able to successfully avoid collision with the human in real-time and reach its goal location in all 10 runs.
@@ VII Conclusion
In this work, we introduced MultiNash-PF, a novel algorithm for efficiently computing multiple local GNEs in CPTGs. By leveraging potential game theory and implicit particle filtering, MultiNash-PF efficiently identifies coarse estimates of local minimizers, which are then refined using optimization solvers to obtain distinct interaction modes. Our numerical simulations demonstrate that MultiNash-PF significantly reduces computation time—by up to 50% compared to baseline methods—while effectively capturing the multimodal nature of multi-agent interactions. Furthermore, our real-world human-robot interaction experiments highlight the algorithm’s ability to reason about multiple interaction modes and resolve conflicts in real-time. These results underscore the potential of MultiNash-PF to enhance interactive motion planning, enabling more adaptive and socially aware autonomous systems.